# Low bit-flip rate probabilistic error cancellation

### v1

@@ VI Conclusion
We have introduced Block-PEC, a variation of the probabilistic error cancellation for error mitigation that takes into account the specifics of the noise model and the propagation of errors to lower the overhead of using noisy quantum computers in lieu of perfect ones. The model that we consider, is that of biased noise qubits as implemented by superconducting cat-qubit architectures.
@@
Our technique provides a provable advantage for circuits where several bias-preserving gates follow one another. It constructs an aggregate noise model that can be inverted all at once instead of a per gate basis. This results in a lower sampling cost for achieving a given precision at the cost of classical pre-computation.
We performed numerical simulations for assessing typical performance on two classes of circuits: first for circuits made of random one and two-qubit bias-preserving gates, and second for specific arrangements of RBS gates. In accordance with our theoretical results, Block-PEC is most efficient for deeper circuits, as the cost of the classical pre-computation increases exponentially with the width of the circuit but only linearly with its depth. For example, for $8$-qubit RBS diagonals (at $p=0.1$) and $8$-qubit RBS pyramids (at $p=0.01$), the overhead is divided by a factor of at least $10$ and $2$ respectively.
Using these simulation results, we argued that for applications such as option pricing or orthogonal neural networks, Block-PEC leads to a reduction of the sampling variance.
@@
We have shown that this advantage carries over whenever sub-circuits made out of several bias-preserving gates can be identified within the full computation. This is possible by performing a hybrid simulation of the full circuit. The bias-preserving sub-circuits are sampled according to the aggregate noise model defined by Block-PEC while the other gates are simply sampled according to the regular PEC technique. Then, one run of the hybrid simulation corresponds to independently sampling each of the sub-circuits and non bias-preserving gates.
@@
This work also exemplifies that cat-qubits architectures can be advantageous to use in the NISQ-regime thanks to their noise model that can be more efficiently handled by error mitigation techniques. As such this raises two related questions: what other error mitigation techniques can benefit from bias noise qubits and, more generally, are there architectures for which error mitigation techniques can be tailored and made particularly efficient?
@@ Appendix A A primer in probabilistic error cancellation
Quantum error mitigation techniques aim to minimize the estimation bias when sampling from noisy circuits. The Probabilistic Error Cancellation (PEC zne-pec ) approach performs a Monte Carlo simulation of the inverse of the noise model, in order to remove this estimation bias: from a generating set of noisy (implementable) gates, the inverted noise model is approximated as a linear combination of noisy gates. Therefore, bounds on the sampling cost of PEC are tied to the quality of available physical implementations.
@@ B.1 Inverting uncorrelated dephasing noise
Typically (see e.g. limits-pec ), uncorrelated dephasing noise channels are defined over a set of qubits $A$ as
@@ B.2 Approximating impure dephasing noise
For cat-qubit architectures, the effective bit flip rate is proportional to $e^{-c\overline{n}}$ where $c$ is a constant and $\overline{n}$ is the mean number of photons (the “cat size”), which comes at the cost of a higher phase flip rate, with a rate increase proportional to $\overline{n}$ QCwithCats .
For a realistic description of the noise model of cat-qubit architectures, one could take the noise channel $\mathsf{L}^{(1)}_{p,q}$ with $q=\frac{e^{c\overline{n}}}{\overline{n}}-1$, so that $\frac{1}{q+1}=\overline{n}e^{-c\overline{n}}$. However, for $c=2$, $q$ is of the order of $10^{3}$ and $10^{7}$ for $\overline{n}$ equal to $5$ and $10$ respectively (which are both reasonable values for $\overline{n}$ under current experiments 100secs-bitflip ; QCwithCats ; reglade2023 ). Therefore, in the context of PEC-like error mitigation protocols, pure dephasing noise channels constitute a good approximation of the noise associated to bias-preserving circuits implemented on cat-qubits.
@@ D.1 Reconfigurable Beam Splitter (RBS) gates
The use of RBS gates for unary data loading has found applications to Orthogonal Neural Networks onn , Nearest Centroid Classification ncc , and solving Partial Differential Equations such as Navier-Stokes, using Quantum Fourier Networks rbs-pde .
RBS gates can also be used to implement the inner product of two vectors $x$ and $y$, by loading the vector $y$ and then unloading (i.e. executing the reversed data loading circuit) the vector $x$. This unary inner product routine is a key element of the implementation of low-depth amplitude estimation low-depth-qae , which finds applications in quantum Monte Carlo simulations Montanaro_2015 , leading notably to an unary quantum algorithm for option pricing unary-option-pricing .
@@ D.2 RBS gates are partially bias-preserving
This property holds not only for $(i)RBS(\theta)$, but for other parity-preserving gates such as $\text{A}(\theta,\phi)$ (see e.g. Barkoutsos_2018 ; Gard_2020 ; vpe-vff ), $\text{Givens}(\theta)$ (see e.g. Anselmetti_2021 ; Kivlichan_2018 ), $\text{XY}(\theta)$ (see e.g. Abrams_2020 ).

### v2

@@ 6 Conclusion
We have introduced Block-PEC, a variation of the probabilistic error cancellation for error mitigation that takes into account the specifics of the noise model and the propagation of errors to lower the overhead of using noisy quantum computers in lieu of perfect ones. The model that we consider corresponds to biased noise qubits as implemented by superconducting cat-qubit architectures.
@@
Our technique provides a provable advantage for single layer circuits and extends to an exponential advantage when several bias-preserving gates follow one another. It constructs an aggregate noise model that can be inverted all at once instead of a per gate basis. This results in a lower sampling cost for achieving a given precision at the cost of classical pre-computation.
We performed numerical simulations for assessing typical performance on several classes of circuits. In accordance with our theoretical results, Block-PEC is most efficient for deeper circuits, as the cost of the classical pre-computation increases exponentially with the width of the circuit but only linearly with its depth. For example, for 9-qubit circuits with $RBS$ gates in three repetitions of a $SWAP$ network, the gain is already about 4 even at $p=0.001$, and for pyramids of $RBS$ gates it is expected to reach 12 for 30 qubits under the same noise strength. Using these simulation results, we argued that for applications such as option pricing or orthogonal neural networks, Block-PEC leads to a reduction of the sampling variance.
@@
We have shown that this advantage carries over whenever sub-circuits made out of several bias-preserving gates can be identified within the full computation. This is possible by performing a hybrid simulation of the full circuit. The bias-preserving sub-circuits are sampled according to the aggregate noise model defined by Block-PEC while the other gates are simply sampled according to the regular PEC technique. Then, one run of the hybrid simulation corresponds to independently sampling each of the sub-circuits and non bias-preserving gates. Similar results could be obtained for Clifford circuits under Pauli noise.
Moreover, applying our per-block reasoning to Noise-Extended PEC (NEPEC) [22], which generalizes both PEC and Zero-Noise Extrapolation (ZNE), means that mitigating errors by block can be an efficient strategy in the context of ZNE.
@@
Finally, this work also exemplifies that cat-qubits architectures can be advantageous to use in the NISQ-regime thanks to their noise model that can be more efficiently handled by error mitigation techniques. As such this raises two related questions: what other error mitigation techniques can benefit from biased noise qubits and, more generally, are there architectures for which error mitigation techniques can be tailored and made particularly efficient for end-to-end applications?
Conflict of Interest.
The authors declare absence of conflict of interest.
@@ Appendix A A primer in probabilistic error cancellation
Quantum error mitigation techniques aim to minimize the estimation bias when sampling from noisy circuits. The Probabilistic Error Cancellation (PEC [32]) approach performs a Monte Carlo simulation of the inverse of the noise model, in order to remove this estimation bias: from a generating set of noisy (implementable) gates, the inverted noise model is approximated as a linear combination of noisy gates. Therefore, bounds on the sampling cost of PEC are tied to the quality of available physical implementations.
@@ B.1 Inverting uncorrelated dephasing noise
Typically (see e.g. [31]), uncorrelated dephasing noise channels are defined over a set of qubits $A$ as
@@ B.2 Approximating impure dephasing noise
For cat-qubit architectures, the effective bit flip rate is proportional to $e^{-c\overline{n}}$ where $c$ is a constant and $\overline{n}$ is the mean number of photons (the “cat size”), which comes at the cost of a higher phase flip rate, with a rate increase proportional to $\overline{n}$ [16].
For a realistic description of the noise model of cat-qubit architectures, one could take the noise channel $\mathsf{L}^{(1)}_{p,q}$ with $q=\frac{e^{c\overline{n}}}{\overline{n}}-1$, so that $\frac{1}{q+1}=\overline{n}e^{-c\overline{n}}$. However, for $c=2$, $q$ is of the order of $10^{3}$ and $10^{7}$ for $\overline{n}$ equal to $5$ and $10$ respectively (which are both reasonable values for $\overline{n}$ under current experiments [5, 16, 26]). Therefore, in the context of PEC-like error mitigation protocols, pure dephasing noise channels constitute a good approximation of the noise associated to bias-preserving circuits implemented on cat-qubits.
@@ D.1 Reconfigurable Beam Splitter (RBS) gates
The use of $RBS$ gates for unary data loading has found applications to Orthogonal Neural Networks [19], Nearest Centroid Classification [18], and solving Partial Differential Equations such as Navier-Stokes, using Quantum Fourier Networks [17].
RBS gates can also be used to implement the inner product of two vectors $x$ and $y$, by loading the vector $y$ and then unloading (i.e. executing the reversed data loading circuit) the vector $x$. This unary inner product routine is a key element of the implementation of low-depth amplitude estimation [15], which finds applications in quantum Monte Carlo simulations [23], leading notably to an unary quantum algorithm for option pricing [25].
@@ D.2 RBS gates are partially bias-preserving
This property holds not only for $(i)RBS(\theta)$, but for other parity-preserving gates such as $\text{A}(\theta,\phi)$ (see e.g. [4, 14, 13]), $\text{Givens}(\theta)$ (see e.g. [2, 20]), $\text{XY}(\theta)$ (see e.g. [1]).